Waste Services Corp (WSC) is looking to open a colossal landfill in Columbia. Examine the possible cash flows below:

Year Cash Flow

0   -29,000

1    11,200

2    13,900

3 15,800

4    12,000

5 -9,400

The company uses an 6.20% discount rate on all of its projects. What is the MIRR of the 'mega dump' using the combination approach?
Answer #1

When you are calculating MIRR using the combination approach you will convert all the cash out flows at time 0 and all cash inflow at the end of the project

Hence, Year 0 cash outflow= -29000-9400/(1+6.2%)^5=$-35,958.33

Year 5 cash inflow= 11200*(1+6.2%)^(5-1)+13900*(1+6.2%)*(5-2)+15800*(1+6.2%)^(5-3)+12000*(1+6.2%)^(5-4)=$89096.09

To calculate MIRR:

0=-35958.33+89096.09/(1+MIRR)^5

or , MIRR= (89096.09/35958.33)^(1/5)-1=19.90%

Hence, MIRR od the mega dump is 19.90%